# Features are fate: a theory of transfer learning in high-dimensional regression

### v1

@@ Appendix A Initialization assumption
To our knowledge, this is the most general assumption on the weight initializations in the literature that leads to the implicit biases that are crucial for our analysis. This initialization scheme generalizes that in Wu et al. (2019); Atanasov et al. (2021).
Appendix B Proofs
B.1 Proof of Theorem 2.2
@@ Appendix B Proofs
Since, $f\in L_{2}(p)$, the sum on right hand side of Eq. (30) converges to some $a<\infty$. We can choose $g=\sqrt{\left|\frac{2\delta e^{\sigma^{2}/2}-a}{a}\right|}\sum_{i=1}^{M}\alpha_{i}\phi_{i}$ which completes the first half of the proof. To prove the result about the KL divergence, can directly calculate $\mathcal{D}_{KL}(p_{f}\|p_{g})$
@@
$\implies\mathcal{L}(t)=\mathcal{L}(0)e^{-2(\alpha^{2}\lambda)^{L-1}t}$ (48)
@@
To prove Lemma B.1 we bound the difference $\lVert\bm{W}_{l}\rVert^{2}_{F}-\lVert\bm{W}_{l}\rVert^{2}_{op}$ which is equal to the norm of the subleading singular vectors of $\bm{W}_{l}$ and show that this bound is proportional to $\alpha^{2}$. The argument here follows that in (Yun et al. (2021)). Taking the trace of both sides in (35) we have
@@
where we have used Lemma B.1 in the third equality to eliminate the inner product between the adjacent singular vectors. The possible factor of $-1$ can be absorbed into the definition of $\bm{u}_{1}$. This implies
@@
B.4 Proof of Theorem 3.5
This proof follows Yun et al. (2021) closely but extends their result to the case $n>d$. We first show that gradient flow converges to a global minimum of the empirical loss (5). We then show that as $\alpha\to 0$, this minimum corresponds to the minimum norm least squares solution.
@@
where $\sigma_{\min}(\bm{W}_{L-1}^{T}\dots\bm{W}_{1}^{T})$ is the smallest singular value of $\bm{W}_{L-1}^{T}\dots\bm{W}_{1}^{T}$. From Eq. (42) - (46) we can bound this quantity away from zero. Then we have
@@
In the case $n>d$, the least squares problem () is overdetermined so the solution is unique. That is, the unique solution is trivially the minimum norm solution. In the case $n\leq d$, there are multiple $\bm{\beta}(t)$ that yield zero training error. Lemma B.1 shows that the parameter matrices are approximately rank one at all times and $\bm{u}_{l+1}$ and $\bm{v}_{l}$ align at the end of training as $\alpha\to 0$, which means that
@@
where we have used the independence of $\bm{\epsilon}$ and $\bm{X}$, as well as the fact that the operator $\bm{X}^{+}\bm{X}$ is the projector onto subspace spanned by the rows of $\bm{X}$, $\bm{P}_{\text{row}(\bm{X})}$. Since the entries of the data matrix $\bm{X}$ are independent Gaussians, the n-dimensional subspace $\text{row}(\bm{X})$ is uniformly random in the Grassmanian manifold $\mathcal{G}_{n,d}$ Vershynin (2018), so $\bm{P}_{\text{row}(\bm{X})}\bm{\beta}_{\mathrm{t}}$ is a random projection of $\bm{\beta}_{\mathrm{t}}$. Then
@@
which is a classic result in the theory of random projections (c.f. Vershynin (2018) Lemma 5.3.2). We now turn to the final term in (95). Let $\{\sigma_{l}\}_{l\leq\min(n,d)}$ be the nonzero singular values of the data matrix $\bm{X}$. Then
@@
In the second line we have introduced the complex variable $z$, which casts the quantity of interest as the $z\to 0$ limit of the normalized expected trace of the resolvent of $\bm{C}$. In the limit of large $n$, this quantity tends to the Stieltjes transform of the Wishart matrix $\mathfrak{g}_{\bm{C}}(z)$, which has a closed form expression (see Potters & Bouchaud (2020) Ch.4 for a proof).
@@
$b$ $=\frac{\beta_{src}^{T}\bm{X}^{T}\bm{y}_{\textrm{t}}}{\bm{\beta}_{\mathrm{s}}^{T}\bm{X}^{T}\bm{X}\bm{\beta}_{\mathrm{s}}}$ (108)
$=\frac{\beta_{src}^{T}\bm{X}^{T}\bm{X}\bm{\beta}_{\mathrm{t}}}{\bm{\beta}_{\mathrm{s}}^{T}\bm{X}^{T}\bm{X}\bm{\beta}_{\mathrm{s}}}+\frac{\beta_{src}^{T}\bm{X}^{T}\bm{\epsilon}}{\bm{\beta}_{\mathrm{s}}^{T}\bm{X}^{T}\bm{X}\bm{\beta}_{\mathrm{s}}}$ (109)
@@
Appendix C ReLU networks
@@
Appendix D Experimental details
D.1 Deep linear models
@@ D.2 ReLU networks
Appendix E Additional Figures
Figure 5: Transferability is not predicted by $\phi$-divergences or integral probability metrics We generate source and target distributions $p_{\mathrm{s}}$, $p_{\mathrm{t}}$ according to the setup in Section 3 and plot the transferability $\mathcal{T}$ (7) as a function of (a) the KL divergence $D_{\mathrm{KL}}(p_{\mathrm{s}}\|p_{\mathrm{t}})$ and (b) the Wasserstein $1$-metric. The KL divergence can be computed exactly in this setting (see Section B.1). $W_{1}$ is computed from finite samples using the algorithm in Sriperumbudur et al. (2009).

### v2

@@ Appendix B Initialization assumption
To our knowledge, this is the most general assumption on weight initializations in the literature that leads to the implicit biases that are crucial for our analysis. This initialization scheme generalizes that in Wu et al. (2019); Atanasov et al. (2021).
Appendix C Fine tuning transferability with finite source data
In the main text, we consider training the source task on the population loss, which can be understood as having infinite data points in the source task. Although the number of source data points typically dominates the number of target data points in real-world applications, it is interesting to consider the scenario in which the dataset sizes are comparable. To this end, we consider a source task with $n_{s}=\gamma_{s}d$ data points and a target task with $n_{t}=\gamma_{t}d$ data points, generalizing the setup of 3. Since the source dataset is now finite, we also add label noise to the source task so that the labels are generated as
$y_{s}$ $=\bm{\beta}_{\mathrm{s}}^{T}\bm{x}+\epsilon_{s}$ (20)
$y_{t}$ $=\bm{\beta}_{\mathrm{t}}^{T}\bm{x}+\epsilon_{t}$ (21)
where $\epsilon_{s}\sim\mathcal{N}(0,\sigma^{2}_{s})$ and $\epsilon_{t}\sim\mathcal{N}(0,\sigma_{t}^{2})$. Aside from these changes, the setup in the following is the same as that in 3. We derive the following expression for the expected generalization error using full fine-tuning, which we prove in Appendix D.9.
Theorem C.1.
Under assumptions above, and assuming the source-target overlap is $\theta$, the expected transferability of the fine-tuned model over the training data is:
$\mathcal{T}_{\mathrm{ft}}=\mathbb{E}_{\mathcal{D}_{t}}\mathcal{R}_{\mathrm{sc}}-\mathbb{E}_{\mathcal{D}_{s},\mathcal{D}_{t}}\mathcal{R}_{\mathrm{ft}}=\begin{cases}\frac{\gamma_{t}-1}{1-\gamma_{s}}\sigma_{s}^{2}\gamma_{s}+\gamma_{s}(\gamma_{t}-1)(1-2\cos\theta)&\gamma_{s},\gamma_{t}\leq 1\\ (\gamma_{t}-1)(1-2\cos\theta)-\frac{\gamma_{t}-1}{1-\gamma_{s}}\sigma_{s}^{2}&\gamma_{s}>1>\gamma_{t}\\ 0&\gamma_{t}>1\end{cases}$ (22)
Note that this expression agrees with (15) in the limit $\gamma_{s}\to\infty$ and that the negative transfer boundary is completely determined by $\gamma_{s}$, since $\gamma_{t}$ only enters the expression via a global multiplicative factor. Secondly, without label noise in the source task, the phase diagram is the same as in Fig 2. That is to say, fine tuning does not depend on the amount of source data if there is no label noise. In the general case, for fixed $\gamma_{s}$, the boundary between positive and negative transferability occurs at a fixed value of $\theta$:
$\cos\theta_{*}=\begin{cases}\frac{1}{2}\left(\frac{\sigma_{s}^{2}+1-\gamma_{s}}{1-\gamma_{s}}\right)&\gamma_{s}<1\\ \frac{1}{2}\left(\frac{\sigma_{s}^{2}+\gamma_{s}-1}{\gamma_{s}-1}\right)&\gamma_{s}>1\\ \end{cases}$ (23)
Therefore, the fine tuning transferability phase diagram with a finite source dataset is qualitatively the same as that with infinite source data (Fig. 2), but the position of the phase boundary is modulated by the size of the source dataset.
Appendix D Proofs
D.1 Proof of Theorem A.2
@@ Appendix D Proofs
Since, $f\in L_{2}(p)$, the sum on right hand side of (31) converges to some $a<\infty$. We can choose $g=\sqrt{\left|\frac{2\delta e^{\sigma^{2}/2}-a}{a}\right|}\sum_{i=1}^{M}\alpha_{i}\phi_{i}$ which completes the first half of the proof. To prove the result about the KL divergence, can directly calculate $\mathcal{D}_{KL}(p_{f}\|p_{g})$
@@
$\implies\mathcal{L}(t)\leq\mathcal{L}(0)e^{-2(\alpha^{2}\lambda)^{L-1}t}$ (49)
@@
To prove Lemma D.1 we bound the difference $\lVert\bm{W}_{l}\rVert^{2}_{F}-\lVert\bm{W}_{l}\rVert^{2}_{op}$ which is equal to the norm of the subleading singular vectors of $\bm{W}_{l}$ and show that this bound is proportional to $\alpha^{2}$. The argument here follows that in ((Yun et al., 2021)). Taking the trace of both sides in (36) we have
@@
where we have used Lemma D.1 in the third equality to eliminate the inner product between the adjacent singular vectors. The possible factor of $-1$ can be absorbed into the definition of $\bm{u}_{1}$. This implies
@@
D.4 Proof of Theorem 3.5
This proof follows (Yun et al., 2021) closely but extends their result to the case $n>d$. We first show that gradient flow converges to a global minimum of the empirical loss (3). We then show that as $\alpha\to 0$, this minimum corresponds to the minimum norm least squares solution.
@@
where $\sigma_{\min}(\bm{W}_{L-1}^{T}\dots\bm{W}_{1}^{T})$ is the smallest singular value of $\bm{W}_{L-1}^{T}\dots\bm{W}_{1}^{T}$. From (43) - (47) we can bound this quantity away from zero. Then we have
@@
In the case $n>d$, the least squares problem () is overdetermined so the solution is unique. That is, the unique solution is trivially the minimum norm solution. In the case $n\leq d$, there are multiple $\bm{\beta}(t)$ that yield zero training error. Lemma D.1 shows that the parameter matrices are approximately rank one at all times and $\bm{u}_{l+1}$ and $\bm{v}_{l}$ align at the end of training as $\alpha\to 0$, which means that
@@
where we have used the independence of $\bm{\epsilon}$ and $\bm{X}$, as well as the fact that the operator $\bm{X}^{+}\bm{X}$ is the projector onto subspace spanned by the rows of $\bm{X}$, $\bm{P}_{\text{row}(\bm{X})}$. Since the entries of the data matrix $\bm{X}$ are independent Gaussians, the n-dimensional subspace $\text{row}(\bm{X})$ is uniformly random in the Grassmanian manifold $\mathcal{G}_{n,d}$ (Vershynin, 2018), so $\bm{P}_{\text{row}(\bm{X})}\bm{\beta}_{\mathrm{t}}$ is a random projection of $\bm{\beta}_{\mathrm{t}}$. Then
@@
which is a classic result in the theory of random projections (c.f. (Vershynin, 2018) Lemma 5.3.2). We now turn to the final term in (96). Let $\{\sigma_{l}\}_{l\leq\min(n,d)}$ be the nonzero singular values of the data matrix $\bm{X}$. Then
@@
In the second line we have introduced the complex variable $z$, which casts the quantity of interest as the $z\to 0$ limit of the normalized expected trace of the resolvent of $\bm{C}$. In the limit of large $n$, this quantity tends to the Stieltjes transform of the Wishart matrix $\mathfrak{g}_{\bm{C}}(z)$, which has a closed form expression (see (Potters & Bouchaud, 2020) Ch.4 for a proof).
@@
$b$ $=\frac{\bm{\beta}_{\mathrm{s}}^{T}\bm{X}^{T}\bm{y}_{\textrm{t}}}{\bm{\beta}_{\mathrm{s}}^{T}\bm{X}^{T}\bm{X}\bm{\beta}_{\mathrm{s}}}$ (109)
$=\frac{\bm{\beta}_{\mathrm{s}}^{T}\bm{X}^{T}\bm{X}\bm{\beta}_{\mathrm{t}}}{\bm{\beta}_{\mathrm{s}}^{T}\bm{X}^{T}\bm{X}\bm{\beta}_{\mathrm{s}}}+\frac{\bm{\beta}_{\mathrm{s}}^{T}\bm{X}^{T}\bm{\epsilon}}{\bm{\beta}_{\mathrm{s}}^{T}\bm{X}^{T}\bm{X}\bm{\beta}_{\mathrm{s}}}$ (110)
@@
D.9 Proof of Theorem C.1
To begin note that by Theorem 3.6, the network will implement the least squares solution to the source task when it is trained with gradient flow on a finite size data set. Denote this vector $\hat{\bm{\beta}}_{\mathrm{bm}}=\bm{X}_{\mathrm{s}}^{+}\bm{y}_{\mathrm{s}}$. By the same reasoning as in Appendix D.8, after fine tuning
$\lim_{\alpha\to 0}\lim_{t\to\infty}\bm{\beta}_{\mathrm{ft}}(t)=\bm{\beta}_{\mathrm{sc}}+(\bm{I}-\bm{P}_{\mathrm{t}})\hat{\bm{\beta}}_{\mathrm{s}}$ (146)
where $\bm{P}_{\mathrm{t}}$ is the orthogonal projector onto the space spanned by the rows of $\bm{X}_{\mathrm{t}}$. With this expression we can write down the expected generalization error of the fine-tuned model
$\mathbb{E}_{\bm{X}_{s},\bm{X}_{t},\bm{\epsilon}_{s},\bm{\epsilon}_{t}}\mathcal{R}_{ft}$ $=\mathbb{E}_{\bm{X}_{s},\bm{X}_{t},\bm{\epsilon}_{s},\bm{\epsilon}_{t}}\lVert\bm{\beta}_{\mathrm{t}}-\bm{\beta}_{\mathrm{sc}}-(\bm{I}-\bm{P}_{\mathrm{t}})\hat{\bm{\beta}}_{\mathrm{s}}\rVert_{2}^{2}$ (147)
$=\mathbb{E}_{\bm{X}_{t},\bm{\epsilon}_{t}}\lVert\bm{\beta}_{\mathrm{t}}-\bm{\beta}_{\mathrm{sc}}\rVert_{2}^{2}+\mathbb{E}_{\bm{X}_{s},\bm{X}_{t},\bm{\epsilon}_{s}}\lVert(\bm{I}-\bm{P}_{\mathrm{t}})\hat{\bm{\beta}}_{\mathrm{s}}\rVert_{2}^{2}-2\mathbb{E}_{\bm{X}_{s},\bm{X}_{t},\bm{\epsilon}_{s}}\bm{\beta}_{\mathrm{t}}^{T}(\bm{I}-\bm{P}_{\mathrm{t}})\hat{\bm{\beta}}_{\mathrm{s}}$ (148)
$=\mathbb{E}_{\mathcal{D}_{t}}\mathcal{R}_{\mathrm{sc}}+\max{(0,1-\gamma_{\mathrm{t}})}\mathbb{E}_{\bm{X}_{s}\bm{\epsilon}_{s}}\lVert\hat{\bm{\beta}}_{\mathrm{s}}\rVert_{2}^{2}-2\mathbb{E}_{\bm{X}_{s},\bm{X}_{t},\bm{\epsilon}_{s}}\bm{\beta}_{\mathrm{t}}^{T}(\bm{I}-\bm{P}_{\mathrm{t}})\hat{\bm{\beta}}_{\mathrm{s}}$ (149)
where we have used the result on random projections employed in the proof of Theorem 3.6 along with the independence of the matrices $\bm{X}_{s}$ and $\bm{X}_{t}$. Let $\bm{P}_{s}$ denote the orthogonal projector onto the rowspace of $\bm{X}_{\mathrm{s}}$. Then we can write
$\mathbb{E}_{\bm{X}_{\mathrm{s}},\bm{\epsilon}_{\mathrm{s}}}\lVert\hat{\bm{\beta}}_{\mathrm{s}}\rVert_{2}^{2}$ $=\mathbb{E}_{\bm{X}_{\mathrm{s}},\bm{\epsilon}_{\mathrm{s}}}\lVert\bm{P}_{\mathrm{s}}\bm{\beta}_{\mathrm{s}}+\bm{X}_{\mathrm{s}}^{+}\bm{\epsilon}_{\mathrm{s}}\rVert_{2}^{2}$ (150)
$=\mathbb{E}_{\bm{X}_{\mathrm{s}},\bm{\epsilon}_{\mathrm{s}}}\lVert\bm{P}_{\mathrm{s}}\bm{\beta}_{\mathrm{s}}\rVert_{2}^{2}+\sigma_{s}^{2}\mathbb{E}_{\bm{X}_{\mathrm{s}}}\mathrm{tr}(\bm{X}_{\mathrm{s}}^{+T}\bm{X}_{\mathrm{s}}^{+})$ (151)
=$=\begin{cases}\gamma_{s}+\frac{\sigma_{s}^{2}\gamma_{s}}{1-\gamma_{s}}&\gamma_{s}<1\\ 1+\frac{\sigma_{s}^{2}}{\gamma_{s}-1}&\gamma_{s}>1\end{cases}$ (152)
where the final line follows from the same reasoning as in the proof of Theorem 3.6 (Appendix D.5). The final term in 149 can be computed using similar techniques to those in Appendix D.8. Let $\bm{\beta}_{\mathrm{t}}=a\bm{\nu}+b\hat{\bm{\beta}}_{\mathrm{s}}$ for some $\bm{\nu}\perp\hat{\bm{\beta}}_{\mathrm{s}}$ in the subspace spanned by $\bm{\beta}_{\mathrm{t}}$ and $\hat{\bm{\beta}}_{\mathrm{s}}$. Then
$\mathbb{E}_{\bm{X}_{s},\bm{X}_{t},\bm{\epsilon}_{s}}\bm{\beta}_{\mathrm{t}}^{T}(\bm{I}-\bm{P}_{\mathrm{t}})\hat{\bm{\beta}}_{\mathrm{s}}$ $=\mathbb{E}_{\bm{X}_{\mathrm{s}},\bm{\epsilon}_{\mathrm{2}}}a\mathbb{E}_{\bm{X}_{\mathrm{t}}}\bm{\nu}^{T}(\bm{I}-\bm{P}_{\mathrm{t}})\hat{\bm{\beta}}_{\mathrm{s}}+\mathbb{E}_{\bm{X}_{\mathrm{s}},\bm{\epsilon}_{\mathrm{2}}}b\mathbb{E}_{\bm{X}_{\mathrm{t}}}\hat{\bm{\beta}}_{\mathrm{s}}^{T}(\bm{I}-\bm{P}_{\mathrm{t}})\hat{\bm{\beta}}_{\mathrm{s}}$ (153)
$=\mathbb{E}_{\bm{X}_{\mathrm{s}},\bm{\epsilon}_{\mathrm{2}}}b\mathbb{E}_{\bm{X}_{\mathrm{t}}}\hat{\bm{\beta}}_{\mathrm{s}}^{T}(\bm{I}-\bm{P}_{\mathrm{t}})\hat{\bm{\beta}}_{\mathrm{s}}$ (154)
$=\mathbb{E}_{\bm{X}_{\mathrm{s}},\bm{\epsilon}_{\mathrm{2}}}b\mathbb{E}_{\bm{X}_{\mathrm{t}}}\lVert(\bm{I}-\bm{P}_{\mathrm{t}})\hat{\bm{\beta}}_{\mathrm{s}}\rVert_{2}^{2}$ (155)
$=\max(0,1-\gamma_{t})\mathbb{E}_{\bm{X}_{\mathrm{s}},\bm{\epsilon}_{\mathrm{2}}}b\lVert\hat{\bm{\beta}}_{\mathrm{s}}\rVert_{2}^{2}$ (156)
$=\max(0,1-\gamma_{t})\mathbb{E}_{\bm{X}_{s}}\bm{\beta}_{\mathrm{t}}\bm{P}_{\mathrm{s}}\bm{\beta}_{\mathrm{s}}$ (157)
$=\cos\theta\max(0,1-\gamma_{t})\min(1,\gamma_{s})$ (158)
In the second line we have used the same reasoning as at the end of Appendix D.8, in the third line we have used the properties of orthogonal projectors, and in the remaining lines we have used properties of random projections that are discussed in detail in Appendix D.5. Putting the results above together completes the proof.
Appendix E ReLU networks
@@
Appendix F Experimental details
F.1 Deep linear models
@@ F.2 ReLU networks
Appendix G Additional Figures
Figure 4: Transferability is not predicted by $\phi$-divergences or integral probability metrics We generate source and target distributions $p_{\mathrm{s}}$, $p_{\mathrm{t}}$ according to the setup in Section 3 and plot the transferability $\mathcal{T}$ (5) as a function of (a) the KL divergence $D_{\mathrm{KL}}(p_{\mathrm{s}}\|p_{\mathrm{t}})$ and (b) the Wasserstein $1$-metric. The KL divergence can be computed exactly in this setting (see Section D.1). $W_{1}$ is computed from finite samples using the algorithm in Sriperumbudur et al. (2009).
Figure 5: Regularizing scratch training eliminates anomalous positive transfer. Simulated linear transfer phase diagram for $L=2$, $\sigma=0.2$, $d=500$ (a) with optimal weight decay in the scratch training and (b) without. To tune the weight decay hyperparameter, we sweep over a grid of $\lambda_{\textrm{wd}}\in[0,10^{-4},10^{-3},10^{-2},10^{-1}]$ and choose the model that has the lowest generalization error. The transfer learning procedure is identical to Fig. 1, only scratch training is altered. In the regularized plot (a), the spike of positive transfer along $\gamma=1$ is eliminated, as the regularized scratch trained model does not undergo double descent.
@@ Appendix G Additional Figures
Figure 9: Regularizing pretrained models toward the lazy regime eliminates negative transfer: We train a two layer ReLU network on the transfer learning task defined by (16) and (17) with $\mu=0.9$, $m=1000$, $m_{*}=100$, $d=100$. During pretraining, we include a regularization term $\lambda\sum_{i=1}^{m}\lVert\bm{w}_{i}-\bm{w}_{i}^{(0)}\rVert_{2}^{2}$ where $\bm{w}_{i}^{(0)}$ is the random initial value of weight vector $\bm{w}_{i}$. This regularization prevents the weights of the network from straying far from their intital values. When $\lambda\to\infty$, features are not updated and model operates in a lazy regime. We generate a sweep of pretrained models for $\lambda\in[0,10^{-4},10^{-3},10^{-2},10^{-1}]$. We then linearly transfer each of these pretrained models to the target task and choose the model with the best generalization error (blue). The transferability degrades with target set size as expected, but the optimally regularized pretrained model avoids negative transfer, while the fully rich model (pink) transfers poorly for nearly all dataset sizes considered.